# FactTrack: Time-Aware World State Tracking in Story Outlines

### v1

@@ 1 Introduction
Figure 1: FactTrack tackles the problems of factual inconsistency and plot redundancy. For factual inconsistency detection, FactTrack tracks a validity interval for each fact to distinguish legitimate contradictions from facts simply changing over time. For plot redundancy detection, our method can represent the timeline in more structured form, making detection easier.
@@
To measure the effectiveness of our approach, we introduce a task for detecting contradictions in the planning procedure, with story outlines serving as our test domain. For event pairs flagged by a method as contradictory, we use GPT-4-Turbo to annotate whether they are actually contradictory, on a 1 to 5 scale. Experimental results show that contradictions flagged by FactTrack (using LLaMA2-7B-Chat as a base LLM) are judged to be real contradictions by a score margin of 0.384 more than a fair baseline using LLaMA2-7B-Chat. Additionally, when FactTrack is run on GPT4, the performance significantly surpasses all baselines, including the one using GPT-4-Turbo as a base model.
@@
    Based on that framework, we develop a method, FactTrack, for detecting time-aware factual contradictions in outlines. FactTrack also serves as a plugin structure for improving factual consistency in text planning and generation.
@@ 2 Related Work
Fact verification is a task widely studied in natural language processing, ranging from verifying scientific claims (Thorne et al., 2018; Wadden et al., 2020) to validating fake news (Wang, 2017; Augenstein et al., 2019). Unlike verifying claims against a database of facts, existing work has also demonstrated the feasibility of performing verification within context (Mihaylova et al., 2019; Shaar et al., 2022). We also draw inspiration from efforts to decompose complex sentences into multiple atomic facts using LLMs (Fan et al., 2020; Kamoi et al., 2023; Min et al., 2023). Compared with prior works, our approach specifically operates on temporal structures, maintaining time-dependent fact validity intervals to handle the dynamic nature of the world state as facts change over time.
@@ 3 Time-Aware Atomic Facts
Existing work focuses on leveraging the semantic decomposition capabilities of LLMs for fine-grained fact-checking (Min et al., 2023). In this session, we propose an event-centric, time-aware atomic fact decomposition method. We track facts on a timeline, building toward the design of our FactTrack system. By analyzing the structure of an event and the representation of the world, we develop a method to better decompose events into atomic facts (Section 3.1) with forward and backward directions. We then discuss how to use knowledge of contradictions between atomic facts to determine the validity intervals of these atomic facts on the timeline (Section 3.2). Through this decomposition of facts and by maintaining validity intervals, our method is particularly effective for time-varying facts as is common in domains such as story writing.
@@ 3.1 Fact Decomposition
In narrative theory, “events” represent transitions in the “world state”, reflecting shifts that impact characters, settings, or the overarching scenario (Hühn et al., 2009). We model these transitions using directional atomic facts, capturing the essence of events as transformations from one state to another. We then employ a novel strategy using LLMs to decompose these events into distinct atomic facts with directions and a time validity interval as shown in Figure 2. These facts are then compared pairwise to assess their interrelations and impacts on the narrative structure. Recognizing that events represent transformations in the world state, we classify these atomic facts into three categories: ‘pre-facts’, ‘post-facts’, and ‘static facts’. pre-facts are those truths that exist before an event takes place, post-facts are truths that emerge following an event, and static facts are those truths that remain unchanged throughout. [^2] Traditional AI methods like STRIPS (Fikes and Nilsson, 1971) discuss preconditions and postconditions. Similar to our pre-facts and post-facts, they model the requirements of the world state before an action and the changes to the world state after an action. We prefer the terms pre-fact and post-fact to better express our meaning of tracking time-varying natural language facts with validity intervals on a timeline. To simplify the problem, we interpret static facts as a straightforward amalgamation of a pre-fact and a post-fact, and henceforth concern ourselves with only the latter two categories. Figure 2 shows an example of how we decompose event $\mathcal{B}$ into pre-facts and post-facts.
@@
The term ‘world state’ corresponds to a set of facts that hold at a particular point in time. Compared with object-based tracking, fact-based tracking provides a more flexible state space. The world state at any given moment represents the maximum set of non-contradicting facts. For example, in a process that moves forward in time, at each new event, the world state is used to cross-check with pre-facts and is updated with post-facts. If a fact in the world state contradicts with a new post-fact, then we drop the former (Figure 2).
@@ 3.2 State Tracking on Timeline
$l_{i}$ $=l+(i-1)\cdot\frac{r-l}{k}$
$r_{i}$ $=l+i\cdot\frac{r-l}{k}$
@@
$l_{1}\leq l_{2}\leq r_{1}\leq r_{2}$
Similar ideas also can be found in Allen’s Interval Algebra (Allen, 1983), Our constraint corresponds to Allen’s Interval Algebra is: $Fact_{1}\ o\ Fact_{2}$, but there is slightly different since the fact interval in our assumption is directional.
Figure 5: The general pipeline for how we maintain our data structure. We begin with a new event (e.g., plot point in a story outline), which we decompose into several pre-facts and post-facts (Decompose Events). For each fact, we determine its validity interval based on the world state (Determine Validity Interval), and then detect any contradictions with existing facts in the world state (Detect Contradictions). If the fact does not contradict any existing fact in the world state, then we update the world state with the new fact (Update World State). Otherwise, we write down details about the contradiction, and rewrite the new event conditioned on the preexisting event and details about the contradiction.
@@ 4.1 World State Maintaining
As shown in the light blue block in Figure 5, we maintain two lists to separately keep track of all pre-facts and post-facts. For each atomic fact, we store its content $s_{i}$, start time $t_{i,begin}$, and end time $t_{i,end}$.
@@ 4.2 Operation Pipeline
    Decompose Events. We decompose a new event into pre-facts and post-facts.
@@
    Determine Validity Interval. For each atomic pre-fact or post-fact, we use the world state to determine its validity interval.
@@
    Detect Contradictions. Based on the validity interval of the current fact, we determine whether the current fact contradicts any existing facts.
@@
    Update World State. If needed, we update existing facts based on the current fact and finally add the current fact to the world state.
Pseudocode is shown in Appendix A.3.
@@ 4.2.1 Decompose Events
In the orange block in Figure 5, we decompose the event into pre-facts, post-facts, and static facts. We use zero-shot prompting with an LLM and parse the output for structure afterward; see details in Appendix A.1. After this step, we execute the following three steps in sequence for each atomic fact.
@@ 4.2.2 Determine Validity Interval
Given a new atomic fact, we now use all the facts in the world state to determine its validity interval, corresponding to the purple block in Figure 5. Taking a post-fact as an example, our initial default validity interval is $[l,\inf)$, where $l$ is the right boundary of the event it came from. We then check the facts in the world state from left to right in order, starting with the ones whose left boundary is greater than $l$, until we find the first post-fact that contradicts with the current fact. We return the left boundary of that post-fact as the right boundary of the current fact; if there is no such contradictory post-fact, we return the default value of $\inf$.
@@ 4.2.3 Detect Contradictions
As shown in the pink block in Figure 5, we retrieve all facts that overlap with the current fact, and check for contradictions. Note that this process can operate forward in time, as shown in Figure 2, as well as backward along the arrow of time. There are multiple ways to define the “overlap” of two directional segments. In our approach, we use the strictest constraint in Figure 4, since we only care about whether there are contradictions on the checkpoints, where the two directional facts begin. By requiring contradiction on both checkpoints, as depicted in Figure 4, our method gains better robustness against errors in fact decomposition or the retrieval process, eliminating many false positives. After this step, if there is a contradiction detected, we can print the contradiction information as shown in the example in Table 2, otherwise we directly update the world state.
@@ 4.3 Contradiction Recognition
We use an NLI model finetuned on outputs annotated by GPT4 (Appendix A.4). By using that model, we can give every pair of facts a contradiction score from 0 to 1 corresponding to how likely they are to contradict each other. If this score is over a set threshold, we say that the two facts contradict in our method. For the update condition and contradiction condition, we use different thresholds since false negatives for the update condition are less harmful than for the contradiction condition—it may not cause any major problems if we miss an update to a fact’s validity interval, but we do not want to ignore an actual contradiction. Additionally, to reduce the computational cost of checking all fact pairs, we also use Contriever (Izacard et al., 2021) as a retrieval model to filter out some irrelevant fact pairs; see filtering details in Appendix A.3.
@@
Full Outline Detection (GPT-4-Turbo) 2.355 $\pm$ 0.163 2.859 $\pm$ 0.149
@@
Table 1: Comparison results of FactTrack (based on LLaMA2-7B-Chat) against corresponding baselines on 90 depth-3 story outlines. The contradiction scores range from 1 to 5, as annotated by GPT-4-Turbo; higher is better. Pairwise Score indicates that the GPT-4-Turbo evaluator only sees the events corresponding to the two facts being checked, while in Context Score, GPT-4-Turbo sees the whole story outline as context. We downsample FactTrack’s detections either randomly (e.g., random 500) or based on top contradiction scores to match the number of detections with baselines for fairer comparison; see Appendices B.3 and C for experiment details. FactTrack based on LLaMA2-7B-Chat achieves performance comparable to GPT4-Turbo while significantly outperforming other baselines. Additionally, when FactTrack is run on GPT4, the performance surpasses all baselines.
Event Pair
Atomic Fact Pairs Detected as Contradict
Table 2: An example pair of events detected by FactTrack as having contradicting atomic facts. The contradiction arises because in event 1.3.3, the group already gains knowledge of hidden secrets that are the reason they have been brought together, which contradicts the indication that before event 3.2, they are unaware about those secrets.
5 Evaluation
5.0.0.0.1 Experiment Setup.
To examine our method’s empirical effectiveness, we apply FactTrack to the task of detecting the contradictions in a story outline. While our method can also work online, detecting problems during the outline generation process and providing feedback, we evaluate only offline in this work. We employ an outline generator similar to the detailed outliner from Yang et al. (2022a), modifying the prompt to include more factual information (Appendix 8). In particular, we follow their paradigm of breadth-first outline expansion (Appendix A.5). To balance length and knowledge density, we use story outlines with three layers, with each event having three sub-events, for 39 events in total. In the evaluation process, we randomly select 90 premises from the WritingPrompts dataset (Fan et al., 2018) and use LLaMA2-7B-Chat (Touvron et al., 2023) to generate the outlines; note that we run FactTrack on LLaMA2-7B-Chat as well. The task can be understood as a detection task: given the outline, we want to retrieve some candidate event pairs, indicating that the model considers them to be contradictory. Outline statistics are shown in Appendix C.
[…]
5.0.0.0.2 Baselines.
Due to the lack of directly applicable prior methods for this task, we designed two baselines as follows.
  1. 1.
    Pairwise Detection on LLaMA2-7B-Chat: We let the large language model view two retrieved event nodes and then make a prediction. This method can be relatively better scaled, but it is more challenging to maintain the temporal validity of facts.
  2. 2.
    Full Outline Detection on GPT series: We let the large language model view the entire text and then use Chain of Thought (CoT) to structurally predict some event nodes that contradict each other. This method is limited by the size of the context window; in our problem, the outline of 39 events along with the prompt length is on average about 4800 tokens, so we run this setting with GPT-3.5-Turbo and GPT-4-Turbo.
Due to the different numbers of positively detected samples, we have downsampled the results, as detailed in Appendix C.
5.0.0.0.3 Metrics.
Due to the difficulty and highly context-dependent nature of annotating event contradictions, we have not found a satisfactory method to clearly label ground truth data, even with human annotations (Appendix 7); on inspection, we found GPT-4-Turbo (OpenAI, 2023) annotations to be higher quality and less noisy. Since the contradiction is not binary and sometimes there is some nuance and context dependency, we score the degree of contradiction in candidate pairs on a scale from 1 to 5 (see Appendix B.3). Using this score annotation method, we check the average score of contradiction in pairs detected by each method, i.e., a higher average score means the method is better at detecting actual contradictions. We designed two variations of this metric based on the context provided to GPT-4-Turbo:
  1. 1.
    Pairwise Score directly labels contradictions just by looking at the given pair of events.
  2. 2.
    Context Score labels the two events in the context of seeing the entire rest of the outline.
5.0.0.0.4 Evaluation Results.
Table 1 shows the results of our experiment. FactTrack on LLaMA2-7B-Chat is significantly better on the two metrics than both Full Outline Detection using GPT-3.5-Turbo and Pairwise Detection on LLaMA2-7B-Chat. We also achieve comparable performance with GPT4-Turbo, despite only using LLaMA2-7B-Chat in our method. When we run FactTrack on GPT4, the performance significantly surpasses all baselines [^3] Note that while we used GPT-4-Turbo to generate the fact contradiction data to finetune our NLI model to work on more complex text, GPT-4-Turbo is not directly used in the pipeline of FactTrack on on LLaMA2-7B-Chat. See details in Appendix A.4. The experimental results confirm our intuition that FactTrack can effectively enhance our ability to detect factual contradictions in story planning, benefiting from the decomposition of events and maintenance of validity intervals of atomic facts.
Tables 2 show a representative example of contradictory event pairs. We can see that our approach performs detailed comparisons at a fine-grained fact level rather than simply checking for contextual continuity. See Appendix E for more examples.
The qualitative inspection also shows that the decomposition step remains a bottleneck for FactTrack, as the nature of language leads to misunderstandings of events before decomposition; LLaMA2-7B-Chat also sometimes generates ambiguous or vague atomic facts. In Appendix F we show a failure case resulting from ambiguity. Moreover, our current method only detects binary contradictions and does not handle more complex situations. Detailed error analysis can be found in Appendix F.
@@ 5 Evaluation
6 Ablation Study
We investigate the relative contribution and alternative choices of the individual modules of FactTrack: Decompose Events, Track Fact (Determine Validity Interval $+$ Update World State) and Detect Contradictions.
@@ 6 Ablation Study
    Without Decompose Events: To understand how our system performs without the Decompose Events module, we replace it with the Chain of Thought (CoT) reasoning method. The CoT prompts are shown in Table 8.
@@
    Without Tracking Facts: In this variant, we remove the Determine Validity Interval and Update World State steps. All facts are considered valid by default.

### v2

@@ 1 Introduction
Figure 1: FactTrack tackles the problems of factual inconsistency and plot redundancy. Note that those problems are based on our observations, to provide a clearer understanding of the problem, with both issues considered together in our pipeline and evaluation. For factual inconsistency detection, FactTrack tracks a validity interval for each fact to distinguish legitimate contradictions from facts simply changing over time. For plot redundancy detection, our method can represent the timeline in more structured form, making detection easier.
@@
To measure the effectiveness of our approach, we introduce a task for detecting contradictions in the planning procedure, with story outlines serving as our test domain. For event pairs flagged by a method as contradictory, we use GPT-4 to annotate whether they are actually contradictory, on a 1 to 5 scale. Experimental results show that contradictions flagged by FactTrack (using LLaMA2-7B-Chat as a base LLM) are judged to be real contradictions by a score margin of 0.384 more than a fair baseline using LLaMA2-7B-Chat. Additionally, when FactTrack is run on GPT4, the performance significantly surpasses all baselines, including the one using GPT-4 as a base model.
@@
    Based on that framework, we develop a method, FactTrack, for detecting time-aware factual contradictions in outlines.
@@ 2 Related Work
Fact verification is a task widely studied in natural language processing, ranging from verifying scientific claims (Thorne et al., 2018; Wadden et al., 2020) to validating fake news (Wang, 2017; Augenstein et al., 2019). Unlike verifying claims against a database of facts, existing work has also demonstrated the feasibility of performing verification within context (Mihaylova et al., 2019; Shaar et al., 2022; Li et al., 2023). We also draw inspiration from efforts to decompose complex sentences into multiple atomic facts using LLMs (Fan et al., 2020; Kamoi et al., 2023; Min et al., 2023). Compared with prior works, our approach specifically operates on temporal structures, maintaining time-dependent fact validity intervals to handle the dynamic nature of the world state as facts change over time.
@@ 3 Time-Aware Atomic Facts
Existing work focuses on leveraging the semantic decomposition capabilities of LLMs for fine-grained fact-checking (Min et al., 2023). In this session, we propose an event-centric, time-aware atomic fact decomposition method. We track facts on a timeline, building toward the design of our FactTrack system. By analyzing the structure of an event and the representation of the world, we develop a method to better decompose events into atomic facts (Section 3.1) with forward and backward directions. We then discuss how to use knowledge of contradictions between atomic facts to determine the validity intervals of these atomic facts on the timeline (Section 3.2). Through this decomposition of facts and by maintaining validity intervals, our method is particularly effective for time-varying facts as is common in domains such as story writing. Examples illustrating the concepts in this section are provided in Appendix A.
@@ 3.1 Fact Decomposition
In narrative theory, events represent transitions in the world state, reflecting shifts that impact characters, settings, or the overarching scenario (Hühn et al., 2009). We model these transitions using directional atomic facts, capturing the essence of events as transformations from one state to another. Atomic fact means a basic and concise statement that conveys a single piece of information (Min et al., 2023), see examples in Figure 2. Then we employ a novel strategy using LLMs to decompose these events into distinct atomic facts with directions and a time validity interval as shown in Figure 2. These facts are then compared pairwise to assess their interrelations and impacts on the narrative structure. Recognizing that events represent transformations in the world state, we classify these atomic facts into three categories: pre-facts, post-facts, and static facts. pre-facts are those truths that exist before an event takes place, post-facts are truths that emerge following an event, and static facts are those truths that remain unchanged throughout. [^2] Traditional AI methods like STRIPS (Fikes and Nilsson, 1971) discuss preconditions and postconditions. Similar to our pre-facts and post-facts, they model the requirements of the world state before an action and the changes to the world state after an action. We prefer the terms pre-fact and post-fact to better express our meaning of tracking time-varying natural language facts with validity intervals on a timeline. To simplify the problem, we interpret static facts as a straightforward amalgamation of a pre-fact and a post-fact, and henceforth concern ourselves with only the latter two categories. Figure 2 shows an example of how we decompose event $\mathcal{B}$ into pre-facts and post-facts.
@@
The term world state corresponds to a set of facts that hold at a particular point in time. Compared with object-based tracking, fact-based tracking provides a more flexible state space. The world state at any given moment represents the maximum set of non-contradicting facts. For example, in a process that moves forward in time, at each new event, the world state is used to cross-check with pre-facts and is updated with post-facts. If a fact in the world state contradicts with a new post-fact, then we drop the former (Figure 2).
@@ 3.2 State Tracking on Timeline
$l_{i}$ $=l+(i-1)\cdot\frac{r-l}{k}$ (1)
$r_{i}$ $=l+i\cdot\frac{r-l}{k}$ (2)
@@
$l_{1}\leq l_{2}\leq r_{1}\leq r_{2}$ (3)
Similar ideas also can be found in Allen’s Interval Algebra (Allen, 1983), Our constraint corresponds to Allen’s Interval Algebra is: $Fact_{1}\ o\ Fact_{2}$, but there is slightly different since because different operations dictate the directionality of facts as we shown in Figure 5.
Figure 5: The general pipeline for how we maintain our data structure. We begin with a new event (e.g., plot point in a story outline), which we decompose into several pre-facts and post-facts (Decompose Events). For each fact, we determine its validity interval based on the world state (Determine Validity Interval), and then detect any contradictions with existing facts in the world state (Detect Contradictions). If the fact does not contradict any existing fact in the world state, then we update the world state with the new fact (Update World State). Otherwise, we write down details about the contradiction, and rewrite the new event conditioned on the preexisting event and details about the contradiction. Note that Determine Validity Interval and Update World State are only between facts in the same direction, while Detect Contradictions are only between facts in different directions.
@@ 4.1 World State Maintaining
As shown in the light blue block in Figure 5, we maintain two lists to keep track of all pre-facts and post-facts. For each atomic fact, we store its content $s_{i}$, start time $t_{i,begin}$, and end time $t_{i,end}$.
@@ 4.2 Operation Pipeline
    Decompose Events. decompose a new event into pre-facts and post-facts.
@@
    Determine Validity Interval. Use the world state to find the validity interval for each atomic pre-fact or post-fact.
@@
    Detect Contradictions. Check if the current fact’s validity interval contradict with existing facts.
@@
    Update World State. If needed, update existing facts as necessary and add the current fact to the world state.
Pseudocode is shown in Appendix B.3.
@@ 4.2.1 Decompose Events
In the orange block in Figure 5, we decompose the event into pre-facts, post-facts, and static facts. We use zero-shot prompting with an LLM and parse the output for structure afterward; see details in Appendix B.1. After this step, we execute the following three steps in sequence for each atomic fact.
@@ 4.2.2 Determine Validity Interval
Given a new atomic fact, we use all the facts in the world state to determine its validity interval (purple block in Figure 5). Taking a post-fact as an example, our initial default validity interval is $[l,\inf)$, where $l$ is the right boundary of the event it came from. We then check the facts in the world state whose left boundary is greater than $l$, in order from left to right, until we find the first contradiction. Then, we set the right boundary of the current fact as the left boundary of the detected fact and return.
@@ 4.2.3 Detect Contradictions
As shown in the pink block in Figure 5, we retrieve all facts that overlap with the current fact, and check for contradictions. Note that this process can operate forward in time, as shown in Figure 2, as well as backward along the arrow of time. There are multiple ways to define the “overlap” of two directional segments. In our approach, we use the strictest constraint in Figure 4, since we only care about whether there are contradictions on the checkpoints, where the two directional facts begin. By requiring contradiction on both checkpoints, as depicted in Figure 4, our method gains better robustness against errors in fact decomposition or the retrieval process, eliminating many false positives. After this step, if there is a contradiction detected, we can print the contradiction information as shown in the examples in Appendix F, otherwise we directly update the world state.
@@ 4.3 Contradiction Recognition
We use an NLI model finetuned on outputs annotated by GPT4 (Appendix B.4). By using that model, we can give every pair of facts a contradiction score from 0 to 1 corresponding to how likely they are to contradict each other. If this score is over a set threshold, we say that the two facts contradict in our method. For the update condition and contradiction condition, we use different thresholds since false negatives for the update condition are less harmful than for the contradiction condition—it may not cause any major problems if we miss an update to a fact’s validity interval, but we do not want to ignore an actual contradiction. Additionally, we also use a retrieval model as a filter to reduce the computational cost, see filtering details in Appendix B.3.
@@
Full Outline Detection (GPT-4) 2.355 $\pm$ 0.163 2.859 $\pm$ 0.149
@@
Table 1: Comparison results of FactTrack (based on LLaMA2-7B-Chat) against corresponding baselines on 90 depth-3 story outlines. The contradiction scores range from 1 to 5, as annotated by GPT-4; higher is better. Pairwise Score indicates that the GPT-4 evaluator only sees the events corresponding to the two facts being checked, while in Context Score, GPT-4 sees the whole story outline as context. We downsample FactTrack’s detections either randomly (e.g., random 500) or based on top contradiction scores to match the number of detections with baselines for fairer comparison; see Appendices C.3 and D for experiment details. FactTrack based on LLaMA2-7B-Chat achieves performance comparable to GPT4-Turbo while significantly outperforming other baselines. Additionally, when FactTrack is run on GPT4, the performance surpasses all baselines.
@@
5 Evaluation
5.0.0.0.1 Experiment Setup.
To examine our method’s empirical effectiveness, we apply FactTrack to the task of detecting the contradictions in a story outline. While our method can also work online, detecting problems during the outline generation process and providing feedback, we evaluate only offline in this work. We employ an outline generator similar to the detailed outliner from Yang et al. (2022a), modifying the prompt to include more factual information (Appendix 8). In particular, we follow their paradigm of breadth-first outline expansion. To balance length and knowledge density, we use story outlines with three layers, with each event having three sub-events, for 39 events in total. We estimate there are 3-5 true contradictions per outline on average (Appendix B.5). In the evaluation process, we randomly select 90 premises from the WritingPrompts dataset (Fan et al., 2018) and use LLaMA2-7B-Chat (Touvron et al., 2023) to generate the outlines; note that we run FactTrack on LLaMA2-7B-Chat as well. The task can be understood as a detection task: given the outline, we want to retrieve some candidate event pairs, indicating that the model considers them to be contradictory. Outline statistics are shown in Appendix D.
[…]
5.0.0.0.2 Baselines.
With no directly applicable prior methods, we designed two baselines:
@@ 5 Evaluation
    Pairwise Detection on LLaMA2-7B-Chat: The model compares two retrieved events to make predictions of contradictions. This approach scales well but struggles with maintaining temporal validity.
@@
    Full Outline Detection on GPT series: The model reviews the entire text using Chain of Thought (CoT) to identify contradictory event nodes. Limited by context window size, this method processes an outline of 39 events under our setting, averaging about 4800 tokens, using GPT-3.5-Turbo and GPT-4.
Due to varying numbers of positive detections, results were downsampled (see Appendix D).
5.0.0.0.3 Metrics.
Due to the complexity and context-dependent nature of annotating event contradictions, we haven’t found a clear method to label ground truth data, even with human input (see Appendix 7). Upon review, GPT-4 (OpenAI, 2023) annotations proved to be higher quality and less noisy. Since contradictions aren’t binary and can be nuanced and depend on context, we score them from 1 to 5 (Appendix C.3). We evaluate methods by the average contradiction score in detected pairs—a higher score indicates better detection. We developed two metric variations using GPT-4 context:
  1. 1.
    Pairwise Score directly labels contradictions by examining pairs of events directly.
  2. 2.
    Context Score labels event pairs within the full outline context.
Classical metrics aren’t applicable without gold labels, but we estimate precision and recall using our metrics in Appendix D.
5.0.0.0.4 Evaluation Results.
Table 1 shows the results of our experiment. FactTrack on LLaMA2-7B-Chat is significantly better on the two metrics than both Full Outline Detection using GPT-3.5-Turbo and Pairwise Detection on LLaMA2-7B-Chat. We also achieve comparable performance with GPT4-Turbo, despite only using LLaMA2-7B-Chat in our method. When we run FactTrack on GPT4, the performance significantly surpasses all baselines [^3] Note that while we used GPT-4 to generate the fact contradiction data to finetune our NLI model to work on more complex text, GPT-4 is not directly used in the pipeline of FactTrack on on LLaMA2-7B-Chat. See details in Appendix B.4. These results confirm that FactTrack effectively enhances contradiction detection in story planning by decomposing events and maintaining validity intervals of atomic facts. Examples are in Appendix F.
Qualitative inspection shows that event decomposition remains a bottleneck, as language ambiguity can lead to misunderstandings before decomposition. Appendix G illustrates a failure case due to ambiguity. Additionally, our method only detects binary contradictions and doesn’t address more complex scenarios. Detailed error analysis is in Appendix G.
5.1 Ablation Study
We examine the roles and alternatives of the modules in FactTrack: Decompose Events, Track Facts (Determine Validity Interval $+$ Update World State), and Detect Contradictions.
  1. 1.
    Without Decompose Events: To assess performance without the Decompose Events module, we substitute it with the Chain of Thought, as detailed in Table 8.
  2. 2.
    Without Track Facts: This variant excludes the Determine Validity Interval and Update World State steps, treating all facts as valid by default.